valkraybff:

The dimensions of a rectangular dining room are 18 feet by 16 feet. If a scale factor of One-fourth is used to make a scale model of the dining room, what is the area of the dining room in the scale model? 18 square feet 36 square feet 48 square feet 72 square feet

Extrinix:

So for scaling, you simply multiply the scale factor by the original length and width. \(\color{lime}{18ft \times 16ft}\) -> Your original area Now if you multiply the scale factor. \(\color{lime}{(18ft \times \dfrac{1}{4}) \times (16ft \times \dfrac{1}{4})}\) -> Your scaled area So simply solve that equation and you have your answer :)

valkraybff:

@extrinix wrote:
So for scaling, you simply multiply the scale factor by the original length and width. \(\color{lime}{18ft \times 16ft}\) -> Your original area Now if you multiply the scale factor. \(\color{lime}{(18ft \times \dfrac{1}{4}) \times (16ft \times \dfrac{1}{4})}\) -> Your scaled area So simply solve that equation and you have your answer :)
WHAT

Extrinix:

It's a pretty simple concept, lol \(\dfrac{1}{4}\) is your scaling factor And since you're making a "model," it's a size decrease (which is why you multiply it)